A number is increased by 15% and then decreased by 10%. Find the net increase or decrease per cent.
1
Expert's answer
2012-12-13T09:52:13-0500
Let's denote the number by N. As the number increased by 15%, then

N1 = N·(100% + 15%)/100% = N·1.15.

As the number decreased by 10%, then

N2 = N1·(100% - 10%)/100% = N1·0.9 = N·0.9·1.15 = N·1.035.

So, the net increase percent is

P = (N2/N) · 100% = (N·1.035/N) · 100% = 103.5%.